1. You are bored so you decide to execute a simple physics experiment. A friend drops an object from above a 1.5m-tall window in your house. Through video analysis, you find that the object travels from the top of the window to the bottom of the window in 0.14 seconds. How high above the top of the window was the object dropped?
1
Expert's answer
2021-10-26T17:19:14-0400
"h=0.5gt^2-H\\h=0.5(9.8)(0.14)^2-1.5=-1.4\ m"

It is 1.4 m below the top of the window.